Please help me solve this equation
what is the linear system?
x-2y=8
x+y=2

Answer by ewatrrr(24785)

 
Hi,
the linear system is the system of the following Linear equations:
x-2y=8
x+y=2
What You are looking for is the solution for this system of Equations
Solving using the Elimination Method
x-2y=8
-x-y=-2 |Multiplying 2nd Eq thru by -1 to eliminate x
-3y = -10
y = 10/3 an x = -4/3 | x+y=2, x = 2-(10/3)
Solution for this System is the ordered pair (-4/3,10/3)
CHECKING our Answer***
x + y = 2
(-4/3)+ 10/3 = 6/3 = 2